OpenStudy (anonymous):

7n+6=4n-9 please show work.

OpenStudy (saifoo.khan):

7n - 4n = -9 - 6 3n = - 15 n = -15/3 n = -5

OpenStudy (anonymous):

n=-5

OpenStudy (saifoo.khan):

Dont copy my answer, jk.

OpenStudy (saifoo.khan):

LOL.